OpenStudy (anonymous):

total amount due on loan using simple interest $700...2%....for 3 years...please help

OpenStudy (anonymous):

742

OpenStudy (anonymous):

I=PRT/100 So, 700x2x3= 4200. 4200/100= 42 Total amount= 700+42= $742.

OpenStudy (anonymous):

ty i figured it out i have one now im not getting

OpenStudy (anonymous):

Post it here.

OpenStudy (anonymous):

simple interest ? and amount due? $600.....5%......9mths

OpenStudy (anonymous):

so two answers the simple interest? and amount due?

OpenStudy (anonymous):

interest is 22.5 amount due is 622.5

OpenStudy (anonymous):

600x5x9=270000. 27000/1200= 22.5 Total amount= 600+22.5=$622.5

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

*600x5x9= 27000

OpenStudy (anonymous):

find the simple interest principle $800...rate 6%....9 months

OpenStudy (anonymous):

Interest= $36 Total amount= $836.

OpenStudy (anonymous):

find the simple interest...principle $2000 Rate 10 and 1/2%....3 and 1/2 years

OpenStudy (anonymous):

Interest= $735. Total amount= $2735.

OpenStudy (anonymous):

21/25 times 10/117 times 27/17= in lowest terms